0.2g of oxygen and 3.17g of a hologen combine separately with the same amount of a metal. What is the equivalent of halogen?

Grade:10

2 Answers

Vikas TU
7 years ago
No. of Eq.s oxygen => 0.2/32 = 1/160
No. of Eq.s. Halogen = 3.17/eq. weight
Now the equivalents should be equal.
then,
equivalent of halogen = wt. of halogen/ no. of equivalents = > 3.17 * 160 = > 507.2g
Krishna
3 years ago
Ans is 127 g 
Since 0.2 grams of oxygen  combines with 3.17 g halogen 
So 8g oxygen combines with 127 g Halogen ......